# Deep Regression Representation Learning with Topology

### v3

@@ Abstract
Most works studying representation learning focus only on classification and neglect regression. Yet, the learning objectives and therefore the representation topologies of the two tasks are fundamentally different: classification targets class separation, leading to disconnected representations, whereas regression requires ordinality with respect to the target, leading to continuous representations. We thus wonder how the effectiveness of a regression representation is influenced by its topology, with evaluation based on the Information Bottleneck (IB) principle.
The IB principle is an important framework that provides principles for learning effective representations. We establish two connections between it and the topology of regression representations. The first connection reveals that a lower intrinsic dimension of the feature space implies a reduced complexity of the representation ${\bf{Z}}$. This complexity can be quantified as the conditional entropy of ${\bf{Z}}$ on the target space ${\bf Y}$, and serves as an upper bound on the generalization error. The second connection suggests learn a feature space that is topologically similar to the target space will better align with the IB principle. Based on these two connections, we introduce PH-Reg, a regularizer specific to regression that matches the intrinsic dimension and topology of the feature space with the target space. Experiments on synthetic and real-world regression tasks demonstrate the benefits of PH-Reg. Code: https://github.com/needylove/PH-Reg.
@@ 1 Introduction
Regression is a fundamental task in machine learning in which input samples are mapped to a continuous target space. Representation learning empowers models to automatically extract, transform, and leverage relevant information from data, leading to improved performance. The information bottleneck (IB) principle (Shwartz-Ziv & Tishby, 2017) provides a theoretical framework and guiding principle for learning effectiveness representations. The IB principle suggests to learn a representation ${\bf{Z}}$ with sufficient information about the target ${\bf Y}$ but minimal information about the input ${\bf X}$. For representation ${\bf{Z}}$, sufficiency keeps all necessary information on ${\bf Y}$, while the minimality reduces ${\bf{Z}}$’s complexity and prevents overfitting. The optimal representation, as specified by Achille & Soatto (2018b, a), is the most useful (sufficient) and minimal. they only specified for classification and neglect regression.
The IB principle is applicable to both classification and regression, in that both learn representations that are minimal and sufficient. However, there are some fundamental differences. For example, classification shortens the distance between features belonging to the same class while elongating the distance between features of different classes; the shortening and elongating of distances can be interpreted as minimality and sufficiency, respectively (Boudiaf et al., 2020). The two effects lead to disconnected representations (Brown et al., 2022a). By contrast, in regression, the representations are shown to be continuous and form an ordinal relationship with respect to the target (Zhang et al., 2023). The disconnected and continuous representations are topologically different, as they have different $0^{th}$ Betti numbers. The $0^{th}$ Betti number represents the connectivity in topology, influencing the ‘shape’ of the feature space [^1] In this work, the feature space represents the set of projected data points, i.e. the manifold, rather than the entire ambient space.. While there are a few works investigating the influence of the representation topology in classification (Hofer et al., 2019; Chen et al., 2019), regression is overlooked. We thus wonder what topology the feature space should have for effective regression and how the topology of the feature space is connected to the IB principle.
@@
  • Based on our connections between the topology of feature space and IB principle, we introduce a regularizer named PH-Reg. Applying PH-Reg achieves significant improvement for coordinate prediction on synthetic datasets and real-world regression tasks super-resolution, age estimation and depth estimation.
@@ 2 Related Works
Topological data analysis. Topological data analysis is a recent field that provides a set of topological and geometric tools to infer robust features for complex data (Chazal & Michel, 2021). It can be coupled with feature learning to ensure that learned representations are robust and reflect the training data’s underlying topology and geometric information (Rieck et al., 2020). It has benefitted diverse tasks ranging from fMRI data analysis (Rieck et al., 2020) to and AI-generated text detection (Tulchinskii et al., 2023). It can also be used as a tool to compare data representations (Barannikov et al., 2021a) and data manifolds (Barannikov et al., 2021b). To learn representations that reflect the topology of the training data, a common strategy is to preserving different dimensional topologically relevant distances of the input space and the feature space (Moor et al., 2020; Trofimov et al., 2023). We follow Moor et al. (2020) to preserve topology information. However, unlike classification, regression’s target space is naturally a metric space rich in topology induced by the metric. Consequently, we leverage the topology of the target space, marking the first exploration of topology specific to effective representation learning for regression.
@@ 3.1 Notations
Consider a dataset $S=\{{\bf x}_{i},{\bf z}_{i},{\bf y}_{i}\}_{i=1}^{N}$ with $N$ samples. sampled from a distribution $P$ with the corresponding label ${\bf y}_{i}\in{\mathcal{Y}}$. To predict ${\bf y}_{i}$, a neural network first encodes the input ${\bf x}_{i}$ to a representation ${\bf z}_{i}\in{\mathbb{R}}^{d}$ before apply a regressor $f$, i.e. $\hat{{\bf y}}_{i}=f({\bf z}_{i})$. The encoder and the regressor $f$ are trained by minimizing a task-specific regression loss ${\mathcal{L}}_{m}$ based on a distance between $\hat{{\bf y}}_{i}$ and ${\bf y}_{i}$, i.e. ${\mathcal{L}}_{m}=g(||\hat{{\bf y}}_{i}-{\bf y}_{i}||_{2})$. Typically, an L2 loss is used, i.e. ${\mathcal{L}}_{m}=\frac{1}{N}\sum_{i}||\hat{{\bf y}}_{i}-{\bf y}_{i}||_{2}$, though more robust variants exist such as L1 or the scale-invariant error (Eigen et al., 2014). Note that the dimensionality of ${\bf y}$ is task-specific and is not limited to 1.We denote ${\bf X},{\bf Y}$, and ${\bf{Z}}$ as random variables representing ${\bf x},{\bf y}$, and ${\bf z}$, respectively.
@@ 3.2 ${\mathcal{I}}{\mathcal{B}}$ purely between ${\bf Y}$ and ${\bf{Z}}$
It is worth mentioning that the assumption given in Theorem 1 holds for most neural networks, as neural networks are commonly deterministic functions. For stochastic representations, we commonly learn a distribution $p({\bf{Z}}|{\bf X})$ approaching a fixed distribution, like the standard Gaussian distribution in VAE. In this case, ${\mathcal{H}}({\bf{Z}}|{\bf X})$ will tend to be a fixed constant for ${\bf{Z}}$. Discussions about the choice of $\beta$, i.e. $\beta\in(0,1)$ or $\beta>1$, and more illustrations are given in Appendix B.
@@ 3.5 Encouraging the Same Intrinsic Dimension
Now, we can establish our first connection, which reveals that ${\mathcal{H}}({\bf{Z}}|{\bf Y})$ is bounded by the ID of the feature space. Note, intrinsic dimension is not a well-defined mathematical object, and different mathematical definitions exist (Ma et al., 2018; Birdal et al., 2021). The manifold is assumed to locally resemble a $n$-dimensional Euclidean space. Intuitively, we can consider the ID as the expectation of $n$ over the distribution of this manifold. We first define Intrinsic Dimension following Ghosh & Motani (2023):
@@ 4 PH-Reg for Regression
${\mathcal{L}}^{\prime}_{d}$ purely encourage the feature space to have a lower intrinsic dimension; sometimes it may even result in an intrinsic dimension lower than that of the target space (see Figure 3, Swiss Roll, where the target space is two-dimensional and the feature space is almost one-dimensional.). In contrast, we wish to lower the ID of the feature space while preventing it from lower than that of the target space. We propose to minimize slope between ${\bf e}=[{\bf e}_{1},{\bf e}_{2},\cdots,{\bf e}_{m}]$ and ${\bf n}$:
@@ 5 Experiments
We compare our method with four methods. 1) Information Dropout (InfDrop) (Achille & Soatto, 2018b). InfDrop serves as an IB baseline. It functions as a regularizer designed based on IB, aiming to learn representations that are minimal, sufficient, and disentangled. 2) Ordinal Entropy (OE) (Zhang et al., 2023). OE acts as a regression baseline. It takes the advantages of classification by learning higher entropy feature space for regression tasks. 3) Birdal’s regularizer (i.e., ${\mathcal{L}}_{d}^{\prime}$) (Birdal et al., 2021) serves as an intrinsic dimension baseline. 4) Topology Autoencoder (i.e., ${\mathcal{L}}_{t}$) (Moor et al., 2020) serves as a topology baseline. Note that the proposed PH-Reg is mainly introduced to verify the established connections, and we do not aim for the state-of-the-art results.
@@ 5.2 Real-World Regression Tasks
We conduct experiments on three real-world regression tasks, including depth estimation (Table 2), super-resolution (Table 3) and age estimation (Table 4). The target spaces of the three tasks are topologically different, i.e. a $1$- dimensional line for depth estimation, $3$-dimensional space for super-resolution and discrete points for age estimation. Results on the three tasks demonstrate that both ${\mathcal{L}}_{t}$ and ${\mathcal{L}}_{d}$ can enhance performance, and combining both further boosts the performance. Detailed settings and more discussions are given in Appendix H.
@@ 5.3 Ablation Studies
ID of different methods: Figure 4(d) displays the intrinsic dimension of the last hidden layer, estimated using TwoNN (Facco et al., 2017), for the testing set of NYU-Depth-v2 from different methods throughout training. While our method is based on the Birdal’s estimator (Birdal et al., 2021), another estimator, TwoNN, captures a decrease in intrinsic dimension when applied ${\mathcal{L}}_{d}$. We observe that without ${\mathcal{L}}_{d}$, the intrinsic dimension tends to increase after epoch $3$, potentially overfitting details, whereas ${\mathcal{L}}_{d}$ prevents such a trend.
@@ 6 Conclusion
7 Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix B Discussions about Theorem 1
In Theorem 2, we consider the case where $\phi$ and $\hat{S}$ are independent, since $z$ in $(x,z,y)\sim P$ is drawn without dependence on the entire $N$ data points $\{x_{i},y_{i}\}_{i=1}^{N}$ in equation (18). Thus, our results are consistent and do not contradict with previous findings. Unlike the previous bounds, our bound is determined by the function $Q$, which characterizes the dispersion or the standard deviation of a distribution by its entropy. The function $Q$ exists for general cases, as the dispersion or the standard deviation and the entropy commonly can be estimated for a specific distribution. We thus can find a function $Q$ to upper bound the relationship on the entropy and its dispersion or the standard deviation.
@@ Appendix E Discussions about the regressor
1. Does the appropriate properties (i.e. lower ID and homeomorphism) depend on the regressor?
@@
Table 6: Depth estimation on NYU-Depth-V2. Here Random represents the encoder is fixed in a random state, while PH-Reg means we first train the encoder purely with PH-Reg for 1 epoch, then fix it and train the regressor.
@@ 2. Topology regularization with “simple” vs highly expressive regressors
For both regressors, our regularizer leads to representations with a higher signal-to-noise ratio (since $\mathcal{H}({\bf{Z}}|{\bf Y})$ and $\mathcal{H}({\bf{Z}}|{\bf Y})$ are minimized). This should make it easier for the regressor to estimate the true underlying signal. However, more expressive regressors have a higher capacity to directly estimate the underlying signal, so the room for improvement from the regularization is reduced. In the extreme case of too much expressiveness, our regularizer may again lead to improvements as it may limit overfitting, by minimizing noise in the learned representation.
@@ Appendix F Preliminaries on Topology
However, the $H_{k}(\text{VR}_{\alpha}({\bf S}))$ is obtained based on a single $\alpha$, which is easily affected by small changes in ${\bf S}$. Thus it is not robust and is of limited use for real-world datasets. The persistent homology considers all the possible $\alpha$ instead of a single one, which results in a sequence of $\beta_{k}$. This is achieved through a nested sequence of simplicial complexes, called filtration: $\text{VR}_{0}({\bf S})\subseteq\text{VR}_{\alpha_{1}}({\bf S})\subseteq\cdots\subseteq\text{VR}_{\alpha_{m}}({\bf S})$ for $0\leq\alpha_{1}\leq\alpha_{m}$. Let $\gamma_{i}=[\alpha_{i},\alpha_{j}]$ be the interval corresponding to a $k$-dimensional hole ‘birth’ at the threshold $\alpha_{i}$ and ‘death’ at the threshold $\alpha_{j}$, we denote $\text{PH}_{k}(\text{VR({\bf S})})=\{\gamma_{i}\}$ the set of ‘birth’ and ‘death’ intervals of the $k$-dimensional holes. We only exploit $\text{PH}_{0}(\text{VR({\bf S})})$ in our PH-Reg, since we exploit the topological autoencoder as the topology part and higher topological features merely increase its runtime. An illustration of the calculation $\text{PH}_{0}(\text{VR({\bf S})})$ is given in Figure  5(b). We define $E({\bf S})=\sum_{\gamma\in\text{PH}_{0}(\text{VR}({\bf S}))}|I(\gamma)|$, where $|I(\gamma)|$ is the length of the interval $\gamma$.

### v4

@@ Abstract
Most works studying representation learning focus only on classification and neglect regression. Yet, the learning objectives and, therefore, the representation topologies of the two tasks are fundamentally different: classification targets class separation, leading to disconnected representations, whereas regression requires ordinality with respect to the target, leading to continuous representations. We thus wonder how the effectiveness of a regression representation is influenced by its topology, with evaluation based on the Information Bottleneck (IB) principle. The IB principle is an important framework that provides principles for learning effective representations. We establish two connections between it and the topology of regression representations. The first connection reveals that a lower intrinsic dimension of the feature space implies a reduced complexity of the representation ${\bf{Z}}$. This complexity can be quantified as the conditional entropy of ${\bf{Z}}$ on the target ${\bf Y}$, and serves as an upper bound on the generalization error. The second connection suggests a feature space that is topologically similar to the target space will better align with the IB principle. Based on these two connections, we introduce PH-Reg, a regularizer specific to regression that matches the intrinsic dimension and topology of the feature space with the target space. Experiments on synthetic and real-world regression tasks demonstrate the benefits of PH-Reg. Code: https://github.com/needylove/PH-Reg.
@@ 1 Introduction
Regression is a fundamental task in machine learning in which input samples are mapped to a continuous target space. Representation learning empowers models to automatically extract, transform, and leverage relevant information from data, leading to improved performance. The information bottleneck (IB) principle (Shwartz-Ziv & Tishby, 2017) provides a theoretical framework and guiding principle for learning effectiveness representations. The IB principle suggests learning a representation ${\bf{Z}}$ with sufficient information about the target ${\bf Y}$ but minimal information about the input ${\bf X}$. For representation ${\bf{Z}}$, sufficiency keeps all necessary information on ${\bf Y}$, while the minimality reduces ${\bf{Z}}$’s complexity and prevents overfitting. The optimal representation, as specified by Achille & Soatto (2018b, a), is the most useful (sufficient) and minimal. Yet, they specified classification and neglect regression.
The IB principle is applicable to both classification and regression in that both learn minimal and sufficient representations. However, there are some fundamental differences. For example, classification shortens the distance between features belonging to the same class while elongating the distance between features of different classes; the shortening and elongating of distances can be interpreted as minimality and sufficiency, respectively (Boudiaf et al., 2020). The two effects lead to disconnected representations (Brown et al., 2022a). By contrast, in regression, the representations are shown to be continuous, connected and form an ordinal relationship with respect to the target (Zhang et al., 2023). The disconnected and connected representations are topologically different, as they have different $0^{th}$ Betti numbers. The $0^{th}$ Betti number represents the connectivity in topology, influencing the ‘shape’ of the feature space [^1] In this work, the feature space represents the set of projected data points, i.e. the manifold, rather than the entire ambient space.. While there are a few works investigating the influence of the representation topology in classification (Hofer et al., 2019; Chen et al., 2019), regression is overlooked. We thus wonder what topology the feature space should have for effective regression and how the topology of the feature space is connected to the IB principle.
@@
  • We introduce a regularizer named PH-Reg based on the established connections. Applying PH-Reg achieves significant improvement in coordinate prediction on synthetic datasets and real-world regression tasks such as super-resolution, age estimation, and depth estimation.
@@ 2 Related Works
Topological data analysis. Topological data analysis is a recent field that provides a set of topological and geometric tools to infer robust features for complex data (Chazal & Michel, 2021). It can be coupled with feature learning to ensure that learned representations are robust and reflect the training data’s underlying topology and geometric information (Rieck et al., 2020). It has benefitted diverse tasks ranging from fMRI data analysis (Rieck et al., 2020) to and AI-generated text detection (Tulchinskii et al., 2023). It can also be used as a tool to compare data representations (Barannikov et al., 2021a) and data manifolds (Barannikov et al., 2021b). To learn representations that reflect the topology of the training data, a common strategy is to preserve different dimensional topologically relevant distances of the input space and the feature space (Moor et al., 2020; Trofimov et al., 2023). We follow Moor et al. (2020) to preserve topology information. However, unlike classification, regression’s target space is naturally a metric space rich in topology induced by the metric, and crucial for the intended task. Consequently, we leverage the topology of the target space, marking the first exploration of topology specific to effective representation learning for regression.
@@ 3.1 Notations
Consider a dataset $S=\{{\bf x}_{i},{\bf z}_{i},{\bf y}_{i}\}_{i=1}^{N}$ with $N$ samples, sampled from a distribution $P$ with the corresponding label ${\bf y}_{i}\in{\mathcal{Y}}$. To predict ${\bf y}_{i}$, a neural network first encodes the input ${\bf x}_{i}$ to a representation ${\bf z}_{i}\in{\mathbb{R}}^{d}$ before apply a regressor $f$, i.e. $\hat{{\bf y}}_{i}=f({\bf z}_{i})$. The encoder and the regressor $f$ are trained by minimizing a task-specific regression loss ${\mathcal{L}}_{m}$ based on a distance between $\hat{{\bf y}}_{i}$ and ${\bf y}_{i}$, i.e. ${\mathcal{L}}_{m}=g(||\hat{{\bf y}}_{i}-{\bf y}_{i}||_{2})$. Typically, an L2 loss is used, i.e. ${\mathcal{L}}_{m}=\frac{1}{N}\sum_{i}||\hat{{\bf y}}_{i}-{\bf y}_{i}||_{2}$, though more robust variants exist such as L1 or the scale-invariant error (Eigen et al., 2014). Note that the dimensionality of ${\bf y}$ is task-specific and is not limited to 1. We denote ${\bf X},{\bf Y}$, and ${\bf{Z}}$ as random variables representing ${\bf x},{\bf y}$, and ${\bf z}$, respectively.
@@ 3.2 ${\mathcal{I}}{\mathcal{B}}$ purely between ${\bf Y}$ and ${\bf{Z}}$
It is worth mentioning that the fixed constant assumption given in Theorem 1 holds for most neural networks, as neural networks are commonly deterministic functions. For stochastic representations, we commonly learn a distribution $p({\bf{Z}}|{\bf X})$ approaching a fixed distribution, like the standard Gaussian distribution in VAE. In this case, ${\mathcal{H}}({\bf{Z}}|{\bf X})$ will tend to be a fixed constant for ${\bf{Z}}$. Discussions about the choice of $\beta$, i.e. $\beta\in(0,1)$ or $\beta>1$, and more illustrations are given in Appendix B.
@@ 3.5 Encouraging the Same Intrinsic Dimension
Now, we can establish our first connection, which reveals that ${\mathcal{H}}({\bf{Z}}|{\bf Y})$ is bounded by the ID of the feature space. Note, intrinsic dimension is not a well-defined mathematical object, and different mathematical definitions exist (Ma et al., 2018; Birdal et al., 2021). We first define Intrinsic Dimension following Ghosh & Motani (2023):
@@
The manifold is assumed to locally resemble a $n$-dimensional Euclidean space. Intuitively, we can consider the ID as the expectation of $n$ over the distribution of this manifold.
@@ 4 PH-Reg for Regression
Intuitively, the growth rate of $E({\bf{Z}}_{n})$ is proportional to the volume of the corresponding manifold; this volume is proportional to the intrinsic dimension. In fact, there is a classical result on the growth rate of (Steele, 1988), showing that the growth rate (i.e. the slop) can constrain the intrinsic dimension.
${\mathcal{L}}^{\prime}_{d}$ purely encourage the feature space to have a lower intrinsic dimension; sometimes it may even result in an intrinsic dimension lower than that of the target space (see Figure 3, Swiss Roll, where the target space is two-dimensional and the feature space is almost one-dimensional.). In contrast, we wish to lower the ID of the feature space while preventing it from being lower than that of the target space. We propose to minimize slope between ${\bf e}=[{\bf e}_{1},{\bf e}_{2},\cdots,{\bf e}_{m}]$ and ${\bf n}$:
@@ 5 Experiments
We compare our method with four methods. 1) Information Dropout (InfDrop) (Achille & Soatto, 2018b). InfDrop serves as an IB baseline. It functions as a regularizer designed based on IB, aiming to learn representations that are minimal, sufficient, and disentangled. 2) Ordinal Entropy (OE) (Zhang et al., 2023). OE acts as a regression baseline. It takes advantage of classification by learning higher entropy feature space for regression tasks. 3) Birdal’s regularizer (i.e., ${\mathcal{L}}_{d}^{\prime}$) (Birdal et al., 2021) serves as an intrinsic dimension baseline. 4) Topology Autoencoder (i.e., ${\mathcal{L}}_{t}$) (Moor et al., 2020) serves as a topology baseline. Note that the proposed PH-Reg is mainly introduced to verify the established connections, and we do not aim for state-of-the-art results.
@@ 5.2 Real-World Regression Tasks
We conduct experiments on three real-world regression tasks, including depth estimation (Table 2), super-resolution (Table 3) and age estimation (Table 4). The target spaces of the three tasks are topologically different, i.e. a $1$- dimensional line for depth estimation, $3$-dimensional space for super-resolution and discrete points for age estimation. Detailed settings, related introductions and more discussions are given in Appendix H.
Results on the three tasks demonstrate that both ${\mathcal{L}}_{t}$ and ${\mathcal{L}}_{d}$ can enhance performance, and combining both further boosts the performance. Specifically, combining both achieves $0.48$ overall improvements (i.e. ALL) on age estimation, a PSNR improvement of $0.096$ on super-resolution for Urban100, and a reduction of $6.7\%$ $\delta_{1}$ error on depth estimation.
@@ 5.3 Ablation Studies
ID of different methods: Figure 4(d) displays the intrinsic dimension of the last hidden layer, estimated using TwoNN (Facco et al., 2017), for the testing set of NYU-Depth-v2 from different methods throughout training. While our method is based on Birdal’s estimator (Birdal et al., 2021), another estimator, TwoNN, captures a decrease in ID when applied ${\mathcal{L}}_{d}$. We observe that without ${\mathcal{L}}_{d}$, the intrinsic dimension tends to increase after epoch $3$, potentially overfitting details, whereas ${\mathcal{L}}_{d}$ prevents such a trend.
@@ 6 Conclusion
Acknowledgement
This research / project is supported by the Ministry of Education, Singapore, under the Academic Research Fund Tier 1 (FY2022).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix B Discussions about Theorem 1
In Theorem 2, we consider the case where $\phi$ and $\hat{S}$ are independent, since $z$ in $(x,z,y)\sim P$ is drawn without dependence on the entire $N$ data points $\{x_{i},y_{i}\}_{i=1}^{N}$ in equation (18). Thus, our results are consistent and do not contradict previous findings. Unlike the previous bounds, our bound is determined by the function $Q$, which characterizes the dispersion or the standard deviation of a distribution by its entropy. The function $Q$ exists for general cases, as the dispersion or the standard deviation and the entropy commonly can be estimated for a specific distribution. We thus can find a function $Q$ to upper bound the relationship on the entropy and its dispersion or the standard deviation.
@@ Appendix E Discussions about the regressor
1. Do the appropriate properties (i.e., lower ID and homeomorphism) depend on the regressor?
@@
Table 6: Depth estimation on NYU-Depth-V2. Here, Random represents the encoder is fixed in a random state, while PH-Reg means we first train the encoder purely with PH-Reg for 1 epoch, then fix it and train the regressor.
@@ 2. Topology regularization with “simple” vs highly expressive regressors
For both regressors, our regularizer leads to representations with a higher signal-to-noise ratio (since $\mathcal{H}({\bf{Z}}|{\bf Y})$ and $\mathcal{H}({\bf{Z}}|{\bf Y})$ are minimized). This should make it easier for the regressor to estimate the true underlying signal. However, more expressive regressors have a higher capacity to estimate the underlying signal directly, so the room for improvement from the regularization is reduced. In the extreme case of too much expressiveness, our regularizer may again lead to improvements as it may limit overfitting, by minimizing noise in the learned representation.
@@ Appendix F Preliminaries on Topology
However, the $H_{k}(\text{VR}_{\alpha}({\bf S}))$ is obtained based on a single $\alpha$, which is easily affected by small changes in ${\bf S}$. Thus, it is not robust and is of limited use for real-world datasets. The persistent homology considers all the possible $\alpha$ instead of a single one, which results in a sequence of $\beta_{k}$. This is achieved through a nested sequence of simplicial complexes, called filtration: $\text{VR}_{0}({\bf S})\subseteq\text{VR}_{\alpha_{1}}({\bf S})\subseteq\cdots\subseteq\text{VR}_{\alpha_{m}}({\bf S})$ for $0\leq\alpha_{1}\leq\alpha_{m}$. Let $\gamma_{i}=[\alpha_{i},\alpha_{j}]$ be the interval corresponding to a $k$-dimensional hole ‘birth’ at the threshold $\alpha_{i}$ and ‘death’ at the threshold $\alpha_{j}$, we denote $\text{PH}_{k}(\text{VR({\bf S})})=\{\gamma_{i}\}$ the set of ‘birth’ and ‘death’ intervals of the $k$-dimensional holes. We only exploit $\text{PH}_{0}(\text{VR({\bf S})})$ in our PH-Reg, since we exploit the topological autoencoder as the topology part and higher topological features merely increase its runtime. An illustration of the calculation $\text{PH}_{0}(\text{VR({\bf S})})$ is given in Figure  5(b). We define $E({\bf S})=\sum_{\gamma\in\text{PH}_{0}(\text{VR}({\bf S}))}|I(\gamma)|$, where $|I(\gamma)|$ is the length of the interval $\gamma$.